# On optimal solutions of classical and sliced Wasserstein GANs with non-Gaussian data

### v1

@@ Appendix E Solution for (22) when $q=d=2$
with the optimal parameters of linear generator $\Theta=\mathbf{U}\Lambda\mathbf{V}^{T}$ where $\Lambda\Lambda^{T}=(\pi\tilde{\sigma}_{x}^{2}/4)\mathbf{I},$ and $\mathbf{U},\mathbf{V}^{T}$ are unitary.
@@
When $d=2$ the target function (53) equals to
$f(\mathbf{S})=\frac{S_{11}+S_{22}}{2}$-
$\frac{2\tilde{\sigma}_{x}}{\Gamma(1/2)}\int^{\infty}_{0}\frac{1}{2}z^{-1/2}\frac{S_{11}(1+zS_{22})+S_{22}(1+zS_{11})}{[(1+zS_{11})(1+zS_{22})]^{3/2}}dz.$ (117)
@@
where change of variable $x=z^{2}$ and $x=\frac{\tan\theta}{\sqrt{D_{11}}}$ are respectively be applied to step (a) and (b). Substituting this results into (117),
@@
It is easy to check (118) is valid besides the first case in (119). Now we prove that (118) is also valid when $S_{11}\neq S_{22}\neq 0$. Using [27, equation 19.4.1 and equation 19.7.2]
$(S_{11}-S_{22})[\frac{\partial f(S_{11},S_{22})}{\partial S_{11}}-\frac{\partial f(S_{11},S_{22})}{\partial S_{22}}]$ (120)
$=$ $\frac{\tilde{\sigma}_{x}}{\Gamma(1/2)\sqrt{S_{22}}}[(S_{11}+S_{22})K(1-\frac{S_{11}}{S_{22}})-2S_{22}E(1-\frac{S_{11}}{S_{22}})].$
@@
is non-negative. For $0\leq m<1$, using [27, equation 19.5.1 and equation 19.5.2],
@@
$\sum_{n=0}^{3}(2n(\frac{1}{2})_{n-1}-m(\frac{1}{2})_{n})\frac{(\frac{1}{2})_{n}}{n!n!}m^{n}$
$=\frac{1}{8}m^{2}+\frac{7}{64}m^{3}+\frac{25}{256}m^{4}>0$
@@
While if $m<0$, using imaginary-modulus transformation of [27, equation 19.7.5] one can transform the problem equivalent to the one with $0\leq m<1$. Specifically,
@@
Hence, (117) is Schur-convex, and thus has global minimum at
@@ Appendix F Estimate of the data CDF $\hat{F}_{\mu}(x)$ in (71)
where the kernel $k_{i}(.)$ can be selected as Gaussian or the one easy to perform CDF inverse. The idea of kernel density estimation is a piecewise density estimation and it converges to the PDF when $|\mathcal{M}|$ is large [30]. To be specific, we first rearrange the data in set $\mathcal{M}$ such that
@@
If $x\geq x_{|\mathcal{M}|}+h$ then $\hat{F}_{\mu}(x)=1$ and if $x\leq x_{1}-h$ then $\hat{F}_{\mu}(x)=0$. Also the inverse CDF

### v2

@@ Appendix E Solution for (23) when $q=d=2$
Now we use elliptic integrals to show the special case of Theorem 2 and (27) when $q=d=2$ as follows.
@@
with the optimal parameters of the linear generator $\Theta=\mathbf{U}\Lambda\mathbf{V}^{T}$ where $\Lambda\Lambda^{T}=(\pi\tilde{\sigma}_{x}^{2}/4)\mathbf{I},$ and $\mathbf{U},\mathbf{V}^{T}$ are unitary.
@@
When $d=2$, the target function (64) equals to
$f(\mathbf{S})=\frac{S_{11}+S_{22}}{2}-\frac{2\tilde{\sigma}_{x}}{\Gamma(1/2)}\int^{\infty}_{0}\frac{1}{2}z^{-1/2}\frac{S_{11}(1+zS_{22})+S_{22}(1+zS_{11})}{[(1+zS_{11})(1+zS_{22})]^{3/2}}dz.$ (119)
@@
where the change of variable $x=z^{2}$ and $x=\frac{\tan\theta}{\sqrt{D_{11}}}$ are respectively applied to step (a) and (b). Substituting this results into (119),
@@
It is easy to check (120) is valid besides the first case in (121). Now we prove that (120) is also valid when $S_{11}\neq S_{22}\neq 0$. Using [26, equation 19.4.1 and equation 19.7.2]
$(S_{11}-S_{22})[\frac{\partial f(S_{11},S_{22})}{\partial S_{11}}-\frac{\partial f(S_{11},S_{22})}{\partial S_{22}}]=\frac{\tilde{\sigma}_{x}}{\Gamma(1/2)\sqrt{S_{22}}}[(S_{11}+S_{22})K(1-\frac{S_{11}}{S_{22}})-2S_{22}E(1-\frac{S_{11}}{S_{22}})].$ (122)
@@
is non-negative. For $0\leq m<1$, using [26, equation 19.5.1 and equation 19.5.2],
@@
$\sum_{n=0}^{3}(2n(\frac{1}{2})_{n-1}-m(\frac{1}{2})_{n})\frac{(\frac{1}{2})_{n}}{n!n!}m^{n}=\frac{1}{8}m^{2}+\frac{7}{64}m^{3}+\frac{25}{256}m^{4}>0$
@@
While if $m<0$, using the imaginary-modulus transformation of [26, equation 19.7.5], one can transform the problem equivalent to the one with $0\leq m<1$. Specifically,
@@
Hence, (119) is Schur-convex, and thus has a global minimum at
@@ Appendix F Estimate of the data CDF $\hat{F}_{\mu}(x)$ in (82)
where the kernel $k_{i}(.)$ can be selected as Gaussian or one easy to perform the CDF inverse. The idea of kernel density estimation is a piecewise density estimation, and it converges to the PDF when $|\mathcal{M}|$ is large [19]. To be specific, we first rearrange the data in set $\mathcal{M}$ such that
@@
If $x\geq x_{|\mathcal{M}|}+h$ then $\hat{F}_{\mu}(x)=1$ and if $x\leq x_{1}-h$ then $\hat{F}_{\mu}(x)=0$. Also, the inverse CDF